# Universal spectra of noisy parameterized quantum circuits

### v1

@@ Abstract
Random unitaries are an important resource for quantum information processing. While their universal properties have been thoroughly analyzed, it is not known what happens to these properties when the unitaries are sampled on the present-day noisy intermediate-scale quantum (NISQ) computers. We implement parameterized circuits, which have been proposed as a means to generate random unitaries, on a transmon platform and model these implementations as quantum maps. To retrieve the maps, a machine-learning assisted tomography is used. We find the spectrum of a map to be either an annulus or a disk depending on the circuit depth and detect an annulus-disk transition. By their spectral properties, the retrieved maps appear to be very similar to a recently introduced ensemble of random maps, for which spectral densities can be analytically evaluated.
@@
What happens to these spectral features when $U^{\mathrm{PR}}_{d}$ are sampled on a present-day quantum computer? As implementations on NISQ [15] circuits will inevitably be non-unitary, analyzing features of the unitary spectra is no longer possible. The ability to address the above question thus hinges on answering the following one: Being in the NISQ realm, where should we look for potentially universal spectral features?
In this Letter we analyze spectra of three- and four-qubit parameterized circuits, implemented on IBM quantum processors [16], by modelling these implementations as completely-positive trace-preserving (CPTP) maps [17]. We find that, depending on the depth of a circuit, the distribution of the eigenvalues of the corresponding map follows the ’Single Ring’ scenario [18, 19], i.e., it comes either in the form of an annulus or a disc. We show that a recently introduced ensemble of random maps [20] provides a framework for a classification of NISQ parameterized circuits using two quantifiers only, the strength of dissipation and its rank.
Figure 1 outlines our retrieval procedure, highlighting its three main components: parameterization, tomography, and optimization. We begin by explaining the procedure and then present the experimental results.
Parameterization. NISQ implementations of quantum circuits can be modeled as completely-positive trace-preserving quantum (CPTP) maps [also referred to as ”channels” [21], ”operations” [22], and ”processes” [23]]. Some recent works have already attempted to approximate NISQ circuits and individual gates [24, 25, 26] using Liouvillians [27], which yield specific types of CPTP maps [28, 29]. Here we consider the most general type of CPTP maps (which we will henceforth refer to as ”maps”), operating in $d=2^{n}$-dimensional Hilbert space, with $n$ being the number of qubits.
@@
Our parameterization is only surjective due to the unitary freedom [21]. From the machine-learning (ML) perspective, the fact that different vectors $\bm{\theta}$ yield the same set of Kraus operators – and therefore the same map – is not disadvantageous as both the search space and the optimal manifold dimensions are increased. —
Since our retrieval procedure is based on tomography data, we need to account for state preparation and measurement (SPAM) errors. We do this by considering a non-ideal initial density operator $\rho_{0}=A_{\rho}A_{\rho}^{\dagger}/\text{Tr}[A_{\rho}A_{\rho}^{\dagger}]$, parameterized by a complex $d\times d$ matrix $A_{\rho}$. As aforementioned, here the surjective nature of the parameterization is potentially advantageous for the optimization. To account for measurements errors, generic positive operator-valued measure (POVM) sets $\{E_{j}\}_{j=0,\cdots,d-1}$ [21], are conventionally used; see, e.g., Refs. [33, 34, 25, 35]. However, in the present-day qubit transmon platforms, classical readout errors are dominant [32], which correspond to diagonally dominant matrices $\{E_{j}\}$, with the magnitude of off-diagonal elements often falling below the resolution threshold. We find that for $n\geqslant 3$ and typical number of shots, $N_{s}\leqslant 2^{11}$, a full-fledged POVM modeling would be highly inefficient and even detrimental [31] to the accuracy of the description of measurement errors as compared to the ’diagonal’ readout description [32]. Therefore, we replace the POVM set with $d\times d$ stochastic corruption matrix $C$, so that $(E_{j})_{kl}=\delta_{kl}C_{jl}$, which is parameterized by a $d\times d$ real matrix $A_{C}$, $C_{jl}=|(A_{\text{C}})_{jl}|/\sum_{k=1}^{d}|(A_{\text{C}})_{kl}|$. In total, SPAM errors are parameterized by real $3d^{2}$ vector $\bm{\omega}=\{\text{Re}(A_{\rho})_{i,j},\text{Im}(A_{\rho})_{k,l},(A_{\text{C}})_{k,l}\}_{k,l=1,..,d}$.
Tomography. Several schemes have been proposed [36, 37, 38, 23] and experimentally realized (for $n=1,2$[39, 40, 41] to retrieve maps from measurements. Many of these schemes are based on the state-map duality (Choi-Jamiołkowski isomorphism) [21], thus reducing the problem to a state tomography; some [40, 41, 23] use the idea of compressed sensing [42] which allows to reduce the number of measurements in the case of low-rank maps.
@@
Optimization. Similar to Ref. [43], we use a quadratic cost function $L(\bm{\theta})=\sum_{\bf{s},\bf{b},j}\left[\hat{p}^{(\bf{s},\bf{b})}_{j}(\bm{\theta})-f^{(\bf{s},\bf{b})}_{j}\right]^{2}$, where $f^{(\bf{s},\bf{b})}_{j}$ is the frequency of the $j$-th binary output obtained in the experiment with Pauli modes $P_{\bf{s}}$ and $P_{\bf{b}}$, and
@@
We benchmark the procedure using different maps, including exponents of Liovillians, and an experiment-relevant SPAM error model [31]. Synthetic data sets are generated by performing $N_{s}=2^{10}$ ’shots’ per measurement mode (as in the experiment). Across all tested models, our procedure consistently retrieves maps with spectral supports closely resembling those of the originals.
Figure 3: Eigenvalues $\{\lambda_{j}\},\leavevmode\nobreak\ j=2,...,d^{2}$ of the maps $\mathcal{T}_{d^{2}}$ retrieved with four-qubit parameterized circuits of varying depth $\ell$ implemented on ibmq_belem. The spectra are invariant under the complex conjugation and only the eigenvalues with $\text{Im}\lambda\geqslant 0$ are plotted. The eigenvalues with $\text{Im}\lambda^{\text{DQ}}\leqslant 0$, of the diluted unitaries $\mathcal{DU}_{d}(p,r)$, obtained as spectral approximations of the maps, are shown together with the corresponding $p$ and $r$ values. Radii of the outer, $R_{+}$ and inner, $R_{-}$, circles are given by $R_{\pm}=\sqrt{(1-p)^{2}\pm p^{2}/r}$.
Results. For the experiment, we adapt one of the parameterized circuits considered in Ref. [13]; see Fig. 2(a). It is composed of $\ell$ blocks, each one consisting of two layers of single-qubit rotations, parameterized with $2n$ angles $\{\varphi_{l},\phi_{l}\}$, $l=1,...,n$, and a ladder of cNOT gates. This particular design allows to quickly reach high expressibility values [13] by gradually increasing $\ell$. For example, when limited to a number of samples $\leqslant 10^{5}$ [31], the expressibility of a four-qubit parameterized circuit of depth $\ell\geqslant 4$ is indistinguishable from that obtained with the Haar-random sampling.
@@
Figure 3 presents the spectra $\{\lambda_{j}\},\leavevmode\nobreak\ j=2,...,d^{2}$ (the leading eigenvalue $\lambda_{1}=1$ is not shown) of the retrieved full-rank maps for random four-qubit parameterized circuits of increasing depths (the results for three-quibit circuits are presented in the SM [31]). The eigenvalues are located inside the unit disk thus highlighting a strong non-unitarity character of the maps. Eigenvalues are arranged in an annulus-like spectral distribution whose outer and inner radii diminishing as $\ell$ increases, eventually undergoing an annulus-disc transition.
@@
To identify ensemble $\mathcal{DU}_{p,r}$ with the spectral support best approximating that of the retrieved map, we minimize the $2$-Wasserstein distance between the Gaussian kernel density estimators of the eigenvalues of the retrieved map and those of a typical member of $\mathcal{DU}_{p,r}$ [31].
@@
Conclusion and outlook. We demonstrated that NISQ parameterized circuits, when modeled as quantum maps, display universal spectral properties. Specifically, their spectral supports, which come in the form of an annulus or a disc, are independent of the particular parameter choices and only depend on the circuit depth.
@@
Supplemental Material for
“Universal spectra of noisy parameterized quantum circuits”
I The parameterized circuit and its expressibility
For the experiment, we adapted a circuit from Ref. [S1], referred therein as ’Circuit 2’. A circuit block, sketched in Fig. 2a in the main text, consists of two layers of single-qubit parametrized rotations and a ladder-like chain of CNOT gates (we replaced $R_{x}$ rotations in the first layer of the original design with $R_{y}$). The block is parameterized with $2n$ angles, $\{\varphi_{1},...,\varphi_{n},\phi_{1},...,\phi_{n}\}$, where $n$ is the number of qubits.
The expressibility [S1] of the circuit $U_{\ell}(\bm{\psi})$, consisting of $\ell$ blocks and parameterized by $2n\ell$-dimensional vector of angles, $\bm{\psi}$, is defined as the Kullback-Leibler (KL) divergence,
$\text{Expr}(N_{\text{sample}})=D_{\text{KL}}\bigl{[}p_{\ell}(F;N_{\text{sample}})|p_{\text{Haar}}(F;N_{\text{sample}})\bigr{]},$ (S1)
between the distribution of the fidelitiy $p_{\ell}(F;N_{\text{sample}})$, $F=\left|\langle\bm{0}|U_{\ell}(\bm{\psi})|\bm{0}\rangle\right|^{2}$, obtained with the circuit by uniformly sampling $N_{\text{sample}}$ times over $2n\ell$ angles, and the distribution corresponding to the Haar random sampling, $p_{\text{Haar}}(F;N_{\text{sample}})$ (note that, in the limit $N_{\text{sample}}\to\infty$, the analytical form of the distribution is known, $p_{\text{Haar}}=(d-1)(1-F)^{d-2}$ [S2]). Finally, as the probability distributions need to be estimated with histograms, there is one more parameter which is the number $N_{\text{bins}}$ of bins per unit interval. We use the same number, $N_{\text{bins}}=75$, as in Ref. [S1].
In Fig. S1, we show the expressibility of the circuit as the function of its depth $\ell$, for two numbers of $N_{\text{sample}}$. Starting from $\ell=4$, the expressibility of the circuit matches the expressiblity of the genuine Haar-random sampling for $N_{\text{sample}}=10^{4}$ and $10^{5}$.
Figure S1: Expressibility, Eq. S1, of the parameterized circuit used in the experiment, as function of the number of concatenated blocks, $\ell$, for two numbers of samples. The number of bins per unit interval is 75.
II POVM set and corruption matrix
@@
$E_{j}=\frac{1}{\sqrt{D}}H_{j}\frac{1}{\sqrt{D}},$ (S2)
@@
By using synthetic benchmarks (as described in the next session), we find that the limitations in the number of measurement modes, $N_{m}$, and shots, $N_{s}$, emerging in the experiment, do not allow us to collect sufficient data to ensure that the retrieved POVMs surpass the corruption matrix in modeling measurement errors. Instead, the POVM approach exhibits inferior performance, making the corruption matrix the preferable option; see Fig. S2.
III Benchmarking
@@ III.1 SPAM Model
$\begin{split}\rho_{in}=c_{1}\left|\bm{0}\right\rangle\left\langle\bm{0}\right|+(1-c_{1})\delta\rho,\\ E^{\text{true}}_{j}=c_{2}\left|\bm{j}\right\rangle\left\langle\bm{j}\right|+(1-c_{2})\delta E_{j}.\end{split}$ (S5)
Here, $c_{1},c_{2}\in[0,1]$ and $\left|\bm{0}\right\rangle\left\langle\bm{0}\right|$ and $\left|\bm{j}\right\rangle\left\langle\bm{j}\right|$ represent the density operator corresponding to the ideal initial state and projectors on the basis states. We mixed these operators, in the convex way, with randomly sampled quantum states $\delta\rho$, and POVM sets $\delta\{E_{j}\}$, respectively. Using them, we calculate the probabilities $\{p_{j}^{\bm{s}}\}$ over binary string outputs $\{\bm{j}\}$, by applying different Pauli strings to the initial state and ’measuring’ the results with the POVMs,
$p_{j}^{\bm{s}}=\mathrm{Tr}[E^{\text{true}}_{j}P_{\bm{s}}\rho_{in}P_{\bm{s}}^{\dagger}],$ (S6)
Finally, we obtain synthetic data, ’frequencies’ $\hat{f}_{j}^{\bm{s}}$, by sampling from the probability distribution, Eq. (S6), $N_{\text{sample}}$ times.
@@
${\text{Fidelity}}(\{E^{\text{true}}_{j}\},\{E_{j}\})=\frac{1}{d}\sum_{i=1}^{d}{\text{Tr}}[\sqrt{E^{\text{true}}_{j}\cdot E_{j}}]\in[0,1].$ (S7)
Figure S2 presents the results of the test.
@@
[…]
Figure S2: POVM Fidelity, Eq. (S7), between ground truth and retrieved SPAM models fitted to synthetic data, as function of the number of shots. Both POVM sets (solid line) and a corruption matrices (dashed line) were optimized.
III.2 Synthetic data from CPTP maps

### v2

# Spectra of noisy parameterized quantum circuits: Single-Ring universality

@@ Abstract
Random unitaries are an important resource for quantum information processing. While their universal properties have been thoroughly analyzed, it is not known what happens to these properties when the unitaries are sampled on the present-day noisy intermediate-scale quantum computers. We implement parameterized circuits, which have been proposed as a means to generate random unitaries, on an IBM Quantum processor and model these implementations as quantum maps. To retrieve the maps, a gradient-based process tomography protocol is developed and used. We find the spectrum of a map to be either an annulus or a disk depending on the circuit depth and detect an annulus-disk transition. By their spectral properties, the retrieved maps appear to be very similar to a recently introduced ensemble of random maps, for which spectral densities can be analytically evaluated. Our results establish, via Dissipative Quantum Chaos theory, a connection between intrinsic properties of present-day noisy intermediate-scale quantum (NISQ) computing platforms and non-Hermitian random matrix theory.
I Introduction
@@
What happens to these spectral features when $U^{\mathrm{PR}}_{d}$ are sampled on a present-day quantum computer? As implementations on a NISQ [15] platform will inevitably be non-unitary, analyzing features of the unitary spectra is no longer possible. The ability to address the above question thus hinges on answering the following one: Being in the NISQ realm, where should we look for potentially universal spectral features?
In this work, we analyze spectra of three- and four-qubit parameterized circuits, implemented on IBM quantum processors [16], by modeling these implementations as completely-positive trace-preserving (CPTP) maps [17]. We find that, depending on the depth of a circuit, the distribution of the eigenvalues of the corresponding map follows the ’Single Ring’ scenario [18, 19], i.e., it comes either in the form of an annulus or a disc. We show that a recently introduced ensemble of random maps [20] provides a framework for a classification of NISQ parameterized circuits using two quantifiers only, the strength of dissipation and its rank. Figure 1 outlines our retrieval procedure, highlighting its three main components: parameterization, tomography, and optimization. We begin by explaining the procedure and then present the experimental results.
II Map parametrization
NISQ implementations of quantum circuits can be modeled as completely-positive trace-preserving quantum (CPTP) maps [also referred to as ”channels” [21], ”operations” [22], and ”processes” [23]]. Some recent works have already attempted to approximate NISQ circuits and individual gates [24, 25, 26] using Liouvillians [27], which yield specific types of CPTP maps [28, 29]. Here we consider the most general type of CPTP maps (which we will henceforth refer to as ”maps”), operating in $d=2^{n}$-dimensional Hilbert space, with $n$ being the number of qubits.
@@ II Map parametrization
Our parameterization is only surjective due to the unitary freedom [21]. From perspective of optimization, the fact that different vectors $\bm{\theta}$ yield the same set of Kraus operators – and therefore the same map – is not disadvantageous as both the search space and the optimal manifold dimensions are increased.
Since our retrieval procedure is based on tomography data, we need to account for state preparation and measurement (SPAM) errors. We do this by considering a non-ideal initial density operator $\rho_{0}=A_{\rho}A_{\rho}^{\dagger}/\text{Tr}[A_{\rho}A_{\rho}^{\dagger}]$, parameterized by a complex $d\times d$ matrix $A_{\rho}$. As aforementioned, here the surjective nature of the parameterization is potentially advantageous for the optimization. To account for measurements errors, generic positive operator-valued measure (POVM) sets $\{E_{j}\}_{j=0,\cdots,d-1}$ [21], are conventionally used; see, e.g., Refs. [31, 32, 25, 33]. However, in the present-day qubit transmon platforms, classical readout errors are dominant [34], which correspond to diagonally dominant matrices $\{E_{j}\}$, with the magnitude of off-diagonal elements often falling below the resolution threshold. We find that for $n\geqslant 3$ and typical number of shots, $N_{s}\leqslant 2^{11}$, a full-fledged POVM modeling would be highly inefficient and even detrimental (see Appendix B) to the accuracy of the description of measurement errors as compared to the ’diagonal’ readout description [34]. Therefore, we replace the POVM set with $d\times d$ stochastic corruption matrix $C$, so that $(E_{j})_{kl}=\delta_{kl}C_{jl}$, which is parameterized by a $d\times d$ real matrix $A_{C}$, $C_{jl}=|(A_{\text{C}})_{jl}|/\sum_{k=1}^{d}|(A_{\text{C}})_{kl}|$. In total, SPAM errors are parameterized by real $3d^{2}$ vector $\bm{\omega}=\{\text{Re}(A_{\rho})_{i,j},\text{Im}(A_{\rho})_{k,l},(A_{\text{C}})_{k,l}\}_{k,l=1,..,d}$.
III Tomography
Several schemes have been proposed [35, 36, 37, 23] and experimentally realized (for $n=1,2$[38, 39, 40] to retrieve maps from measurements. Many of these schemes are based on the state-map duality (Choi-Jamiołkowski isomorphism) [21], thus reducing the problem to a state tomography; some [39, 40, 23] use the idea of compressed sensing [41] which allows to reduce the number of measurements in the case of low-rank maps.
@@ III Tomography
IV Optimization
Similar to Ref. [42], we use a quadratic cost function $L(\bm{\theta})=\sum_{\bf{s},\bf{b},j}\left[\hat{p}^{(\bf{s},\bf{b})}_{j}(\bm{\theta})-f^{(\bf{s},\bf{b})}_{j}\right]^{2}$, where $f^{(\bf{s},\bf{b})}_{j}$ is the frequency of the $j$-th binary output obtained in the experiment with Pauli modes $P_{\bf{s}}$ and $P_{\bf{b}}$, and
@@ IV Optimization
We benchmark the procedure using different maps, including exponents of Liovillians, and an experiment-relevant SPAM error model (see Appendix C). Synthetic data sets are generated by performing $N_{s}=2^{10}$ ’shots’ per measurement mode (as in the experiment). Across all tested models, our procedure consistently retrieves maps with spectral supports closely resembling those of the originals.
Figure 3: Eigenvalues $\{\lambda_{j}\},\leavevmode\nobreak\ j=2,...,d^{2}$ of the maps $\mathcal{T}_{d^{2}}$ retrieved with four-qubit parameterized circuits of varying depth $\ell$ implemented on ibmq_belem. The spectra are invariant under the complex conjugation and only the eigenvalues with $\text{Im}\lambda\geqslant 0$ are plotted. The eigenvalues with $\text{Im}\lambda^{\text{DQ}}\leqslant 0$, of the diluted unitaries $\mathcal{DU}_{d}(p,r)$, obtained as spectral approximations of the retieved maps, are shown together with the corresponding $p$ and $r$ values. Radii of the outer, $R_{+}$, and inner, $R_{-}$, circles are given by $R_{\pm}=\sqrt{(1-p)^{2}\pm p^{2}/r}$. For $\ell=40$, the eigenvalues of the corresponding diluted unitaries filled a disc.
V Results
For the experiment, we adapt one of the parameterized circuits considered in Ref. [13]; see Fig. 2(a). It is composed of $\ell$ blocks, each one consisting of two layers of single-qubit rotations, parameterized with $2n$ angles $\{\varphi_{l},\phi_{l}\}$, $l=1,...,n$, and a ladder of cNOT gates. This particular design allows to quickly reach high expressibility values [13] by gradually increasing $\ell$. For example, when limited to a number of samples $\leqslant 10^{5}$, the expressibility of a four-qubit parameterized circuit of depth $\ell\geqslant 4$ is indistinguishable from that obtained with the Haar-random sampling (see Appendix A).
@@ V Results
Figure 3 presents the spectra $\{\lambda_{j}\},\leavevmode\nobreak\ j=2,...,d^{2}$ (the leading eigenvalue $\lambda_{1}=1$ is not shown) of the retrieved full-rank maps for random four-qubit parameterized circuits of increasing depths (the results for three-quibit circuits are presented in Appendix D). The eigenvalues are located inside the unit disk thus highlighting a strong non-unitarity character of the maps. Eigenvalues are arranged in an annulus-like spectral distribution whose outer and inner radii diminishing as $\ell$ increases, eventually undergoing an annulus-disc transition.
@@
To identify ensemble $\mathcal{DU}_{p,r}$ with the spectral support best approximating that of the retrieved map, we minimize the $2$-Wasserstein distance between the Gaussian kernel density estimators of the eigenvalues of the retrieved map and those of a typical member of $\mathcal{DU}_{p,r}$ (see Appendix C).
@@
VI Conclusion
We demonstrated that NISQ parameterized circuits, when modeled as quantum maps, display universal spectral properties. Specifically, their spectral supports, which come in the form of an annulus or a disc, are independent of the particular parameter choices and only depend on the circuit depth.
@@ VII Acknowledgements
Appendix A The parameterized circuit and its expressibility
For the experiment, we adapted a circuit from Ref. [13], referred therein as ’Circuit 2’. A circuit block, sketched in Fig. 2a in the main text, consists of two layers of single-qubit parametrized rotations and a ladder-like chain of CNOT gates (we replaced $R_{x}$ rotations in the first layer of the original design with $R_{y}$). The block is parameterized with $2n$ angles, $\{\varphi_{1},...,\varphi_{n},\phi_{1},...,\phi_{n}\}$, where $n$ is the number of qubits.
The expressibility [13] of the circuit $U_{\ell}(\bm{\psi})$, consisting of $\ell$ blocks and parameterized by $2n\ell$-dimensional vector of angles, $\bm{\psi}$, is defined as the Kullback-Leibler (KL) divergence,
$\text{Expr}(N_{\text{sample}})=D_{\text{KL}}\bigl{[}p_{\ell}(F;N_{\text{sample}})|p_{\text{Haar}}(F;N_{\text{sample}})\bigr{]},$ (4)
between the distribution of the fidelitiy $p_{\ell}(F;N_{\text{sample}})$, $F=\left|\langle\bm{0}|U_{\ell}(\bm{\psi})|\bm{0}\rangle\right|^{2}$, obtained with the circuit by uniformly sampling $N_{\text{sample}}$ times over $2n\ell$ angles, and the distribution corresponding to the Haar random sampling, $p_{\text{Haar}}(F;N_{\text{sample}})$ (note that, in the limit $N_{\text{sample}}\to\infty$, the analytical form of the distribution is known, $p_{\text{Haar}}=(d-1)(1-F)^{d-2}$ [S2]). Finally, as the probability distributions need to be estimated with histograms, there is one more parameter which is the number $N_{\text{bins}}$ of bins per unit interval. We use the same number, $N_{\text{bins}}=75$, as in Ref. [13].
In Fig. 5, we show the expressibility of the circuit as the function of its depth $\ell$, for two numbers of $N_{\text{sample}}$. Starting from $\ell=4$, the expressibility of the circuit matches the expressiblity of the genuine Haar-random sampling for $N_{\text{sample}}=10^{4}$ and $10^{5}$.
Figure 5: Expressibility, Eq. 4, of the parameterized circuit used in the experiment, as function of the number of concatenated blocks, $\ell$, for two numbers of samples. The number of bins per unit interval is 75.
Appendix B POVM set and corruption matrix
@@
$E_{j}=\frac{1}{\sqrt{D}}H_{j}\frac{1}{\sqrt{D}},$ (5)
@@
By using synthetic benchmarks (as described in the next session), we find that the limitations in the number of measurement modes, $N_{m}$, and shots, $N_{s}$, emerging in the experiment, do not allow us to collect sufficient data to ensure that the retrieved POVMs surpass the corruption matrix in modeling measurement errors. Instead, the POVM approach exhibits inferior performance, making the corruption matrix the preferable option; see Appendix B.
Appendix C Benchmarking
@@ C.1 SPAM Model
$\begin{split}\rho_{in}&=c_{1}\left|\bm{0}\right\rangle\left\langle\bm{0}\right|+(1-c_{1})\,\delta\rho,\\ E^{\text{true}}_{j}&=c_{2}\left|\bm{j}\right\rangle\left\langle\bm{j}\right|+(1-c_{2})\,\delta E_{j}.\end{split}$ (8)
Here, $c_{1},c_{2}\in[0,1]$ and $\left|\bm{0}\right\rangle\left\langle\bm{0}\right|$ and $\left|\bm{j}\right\rangle\left\langle\bm{j}\right|$ represent the density operator corresponding to the ideal initial state and projectors on the basis states. We mixed these operators, in the convex way, with randomly sampled quantum states $\delta\rho$ and POVM elements $\delta E_{j}$, respectively. Using them, we calculate the probabilities $\{p_{j}^{\bm{s}}\}$ over binary string outputs $\{\bm{j}\}$ by applying different Pauli strings to the initial state and “measuring” the results with the POVMs,
$p_{j}^{\bm{s}}=\mathrm{Tr}\bigl{[}E^{\text{true}}_{j}\,P_{\bm{s}}\,\rho_{in}\,P_{\bm{s}}^{\dagger}\bigr{]}.$ (9)
Finally, we obtain synthetic data—“frequencies” $\hat{f}_{j}^{\bm{s}}$—by sampling from the distribution in Eq. (9) $N_{\text{sample}}$ times.
Figure 6: POVM Fidelity, Eq. (10), between ground truth and retrieved SPAM models fitted to synthetic data, as a function of the number of shots. Both POVM sets (solid line) and corruption matrices (dashed line) were optimized.
Figure 7: Spectra of original maps (blue dots), Eqs. (8),(9), and of the retrieved maps (orange dots) for $n=3$ (upper panel) and $n=4$ (lower panel). Parameters of the map are (a) $\alpha=1$, $\beta=0.1$, $r=1$, (b) $\alpha=10^{4}$, $\beta=10^{-3}$, $r=16$, (c) $\alpha=10^{2}$, $\beta=10^{-3}$, $r=16$, (d) $\alpha=1$, $\beta=10^{-2}$, $r=8$.
@@
${\text{Fidelity}}(\{E^{\text{true}}_{j}\},\{E_{j}\})=\frac{1}{d}\sum_{i=1}^{d}{\text{Tr}}[\sqrt{E^{\text{true}}_{j}\cdot E_{j}}]\in[0,1].$ (10)
Figure 6 presents the results of the test.
@@
C.2 Synthetic data from CPTP maps